OpenStudy (anonymous):

FAN AND MEDAL PLZ HELP which pair of ratios forms a proportion? A 4/5 and 16/24 B 9/15 and 3/4 C 5/8 and 30/48 D 9/10 and 40/45

OpenStudy (anonymous):

NEED HELP

OpenStudy (anonymous):

C

OpenStudy (dayakar):

a/b = c/d [a,b,c,d are in proportion

OpenStudy (dayakar):

5/8------(1) 30/48 = 5*6/8*6 = 5/8-------(2) (1) =(2)

OpenStudy (anonymous):

MEDAL PLZ